The manager of a movie theater found that Staurday's sales were $3675. He knew that a total of 650 tickets were sold Satueday. Adult tickets cost $7.50, and children's tickets cost $4.50.
1. Write equtions to represent the number of tickets sold and the amount of money collected.
2. How many of each kind if ticket were sold?

Answer by checkley77(12844)

7.5x+4.5(650-x)=3675
7.5x+2925-4.5x=3675
3x=3675-2925
3x=750
x=750/3
x=250 adult tickets were sold.
650-250=400 children tickets were sold.
proof:
7.5*250+4.5*400=3675
1875+1800=3675
3675=3675